The longest distance of the point (a,0) from the curve 2x2 + y2 - 2x = 0 is given by

Grade:12th pass

1 Answers

Aakash
8 years ago
sorry curve is 2x^{2}+y^{2} -2x = 0
and options are
(A) \sqrt{1+2a+2a^{2}}                               (B) \sqrt{1+2a-a^{2}}
(C) \sqrt{1-2a+2a^{2}}                               (D) \sqrt{1-2a+a^{2}}